OpenStudy (anonymous):

Which is the correct solution to the expression 3 + 5 ^ 2? Use the scientific calculator given in this course.

OpenStudy (prettygirl_shynice):

want me to help

OpenStudy (anonymous):

first square 5 which is?

OpenStudy (jaynees):

3 + 5^2 = 3 + 25 = 28

OpenStudy (prettygirl_shynice):

okay frist

OpenStudy (anonymous):

i got it will you guys help with the rest of the test

OpenStudy (anonymous):

@JayneES no direct answers. it's against the code of conduction

OpenStudy (jaynees):

I'm new I wasn't aware

OpenStudy (prettygirl_shynice):

1. 3+5²

OpenStudy (anonymous):

Each of the four sides of a swimmimg pool measures 9 meters. The pool is 5 meters deep. How much water will be needed to fill it completely?

OpenStudy (anonymous):

help?

OpenStudy (jaynees):

This is a volume question, Rose_Marie2016

OpenStudy (prettygirl_shynice):

so then you stay to yourself what 3+25 and that will be your final answer

OpenStudy (prettygirl_shynice):

got it

OpenStudy (anonymous):

\[3 + 5^{2}\] => 3 + 25 = 28

OpenStudy (prettygirl_shynice):

@Rose_Marie2016 got it

OpenStudy (anonymous):

yes thank you @Prettygirl_shynice